### v1

@@ 5 Conclusion
The implementation of Quantum ResNet demonstrates a promising avenue towards creating complex QML models that are resistant to classical simulation, and at the same time are capable of avoiding the trainability issues associated with barren plateaus. Although we highlight this potential in Section 2.4 it is still a question of further research as to finding bounds on this claim and characterising the cases in which this may or may not be useful in general. We also demonstrated that by parameterising the strength of the residual connections through $\beta_{l}$ terms, one can arbitrarily increase the lower bound probability of success of the LCU method. This provides an avenue to tackle the key problem of the LCU method, which is that the implementations can only be performed probabilistically.
We also demonstrated an implementation of quantum native average pooling layers commonly applied in convolutional neural networks. These layers have demonstrated success in classical techniques and therefore have the potential to improve quantum convolutional neural networks. Furthermore, we showed that given a quantum encoded image, the average pooling layers can be implemented without dependency on the size of the image due to the parallelism inherent in quantum circuits, offering potential computational speedups for large images.
Finally, we demonstrated an integration of irreducible representation projections into a quantum framework, allowing for the encoding of data symmetries directly into the quantum state, which has the potential to enhance the model’s generalisation capabilities across various data structures. This general irreducible subspace projection framework includes previous work regarding permutation invariant encodings for point clouds heredge2023permutation as a special case, while in Section 4.3.2 we introduced a novel rotationally invariant encoding for point cloud data using the $S_{4}$ group. We further introduced in Section 4.4 a method of parameterising the amount of symmetry in an encoding as a tuneable hyperparameter. This added flexibility resulted in an improved performance for point cloud data when encoding an intermediate amount of permutation symmetry when compared to either a fully permutation invariant or noninvariant encoding. These were illustrative examples of how the framework could be utilised and there could be many possible symmetries and datasets for which this framework could be similarly adapted and optimised.
@@
J. Heredge envisaged the project, initially focused on irreducible subspace projections, with M. Sevior and L. Hollenberg as an extension of their previous work heredge2023permutation . J. Heredge devised the ResNet and Average Pooling layers applications. The theorems and numerical simulations were developed by J. Heredge with the help of M. West. M. Sevior and L. Hollenberg supervised the overall project. All authors contributed in reviewing the manuscript.
@@ Appendix A Conjugacy Class Implementation
In the implementation of irreducible subspace projections detailed in Section 4, an operator $\hat{\chi}$ was required that contained the character of every element in the group $G$. This meant that the dimension of the matrix scaled with $\rvert G\rvert$, which can be problematic for some groups, such as the permutation group $S_{n}$ where $\rvert S_{n}\rvert=n!$. To tackle this, in this section we will consider a unitary operator $\tilde{\chi}$ that contains the character for each conjugacy class in the group, with additional ancilla qubit registers per conjugacy class prepared in an equal superposition that then acts to distribute these characters to the corresponding group elements. This effectively only requires encoding the character table of the group into a unitary operator, and hence $\tilde{\chi}$ scales with the number of conjugacy classes of $G$. This has the potential to drastically reduce the size of the unitary that must be applied, at the expense of requiring additional ancilla qubits. We show that utilising this circuit structure recreates the result of Theorem 4 and is therefore a valid alternative approach.
The character orthogonality theorem Huppert1998 stated previously in Equation 86 relates to a sum over all the elements of the group. As the character is the same for all elements in a conjugacy class, this can be adjusted to a sum over all conjugacy classes $\nu_{i}\in\mathcal{C}$ by introducing a term $d_{\nu_{i}}=\rvert\nu_{i}\rvert$ that represents the number of group elements in the conjugacy class $\nu_{i}$. Therefore, the character orthogonality theorem can be written
$\frac{1}{\rvert G\rvert}\sum_{{\nu_{i}}\in\mathcal{C}}d_{\nu_{i}}\chi_{j}({\nu_{i}})\chi_{k}({\nu_{i}})^{*}=\delta_{j,k}.$ (112)
@@
$\tilde{\chi}\lvert r\rangle=\frac{1}{\sqrt{\rvert G\rvert}}\sum_{{\nu_{i}}\in\mathcal{C}}\sqrt{d_{\nu_{i}}}\chi_{r}({\nu_{i}})^{*}\lvert{\nu_{i}}\rangle,$ (113)
@@
This successfully allows for the construction of a unitary matrix $\tilde{\chi}$ that scales in size with the number of conjugacy classes as opposed to the number of group elements. We therefore have satisfied the constraint that $\tilde{\chi}$ is unitary and can be implemented on a quantum device.
@@
where $\lvert l_{\nu_{i}}\rangle$ are basis states $\{\lvert l_{\nu_{i}}\rangle\}\in(\mathbb{C}^{2})^{\otimes k_{\nu_{i}}}$ of the $2^{k_{\nu_{i}}}$ dimensional Hilbert space for the $k_{\nu_{i}}$ qubit ancilla register $\lvert\omega_{\nu_{i}}\rangle$ that is encoded on $k_{\nu_{i}}$ qubits, which can be taken to be the computational basis states. The groups elements $l_{\nu_{i}}$ that are contained in the conjugacy class $\nu_{i}$ now label the basis states in register $\lvert\omega_{\nu_{i}}\rangle$. Each register has the requirement $k_{\nu_{i}}\geq\log(d_{\nu_{i}})$ to ensure that there are sufficient qubits in the register such that each element in ${\nu_{i}}$ has a corresponding basis state $\lvert l_{\nu_{i}}\rangle$.
@@
$\otimes\sum_{r=1}^{R}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}\frac{a_{r}n_{r}}{\sqrt{d_{\nu_{i}}}}\chi_{r}({\nu_{i}})^{*}\lvert l_{\nu_{i}}\rangle\lvert{\nu_{i}}\rangle U_{{\nu_{i}},l_{\nu_{i}}}\lvert\psi\rangle\Big{)}$
$=\sum_{\nu_{i}}\bigotimes_{{\nu_{i}}^{\prime}\neq{\nu_{i}}}\lvert\omega_{{\nu_{i}}^{\prime}}\rangle\sum_{r=1}^{R}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}\frac{a_{r}n_{r}\chi_{r}({\nu_{i}})^{*}}{\Omega\sqrt{d_{\nu_{i}}\rvert G\rvert}}\lvert l_{\nu_{i}}\rangle\lvert{\nu_{i}}\rangle$
@@
where we are being flexible with the precise positioning in the tensor product of the unused registers $\lvert\omega_{{\nu_{i}}^{\prime}\neq{\nu_{i}}}\rangle$ for each ${\nu_{i}}^{\prime}$ term for the sake of readability. The registers $\lvert\omega_{{\nu_{i}}}\rangle$ that contain the states $\lvert l_{\nu_{i}}\rangle$ will in fact no longer be used in the circuit and can be ignored. They have performed their role in ensuring that the correct $U_{{\nu_{i}},l_{\nu_{i}}}$ was applied to the state $\lvert\psi\rangle$. This means we can simplify the expression to
$=\sum_{\nu_{i}}\sum_{r=1}^{R}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}\frac{a_{r}n_{r}\chi_{r}({\nu_{i}})^{*}}{\Omega\sqrt{d_{\nu_{i}}\rvert G\rvert}}\lvert{\nu_{i}}\rangle U_{{\nu_{i}},l_{\nu_{i}}}\lvert\psi\rangle.$ (121)
@@
$\langle b_{1}\rvert\tilde{\chi}=\frac{1}{\sqrt{\rvert G\rvert}}\sum_{{\nu_{i}}\in\mathcal{C}}\sqrt{d_{\nu_{i}}}\chi_{1}({\nu_{i}})\langle{\nu_{i}}\rvert.$ (122)
@@
$\Big{(}\langle b_{1}\rvert\tilde{\chi}\Big{)}\cdot\Big{(}\hat{S}\big{(}\bigotimes_{{\nu_{i}}^{\prime}}\lvert\omega_{{\nu_{i}}^{\prime}}\rangle\otimes\tilde{\chi}\hat{\gamma}\lvert b_{1}\rangle\otimes\lvert\psi\rangle\big{)}\Big{)}$
@@ Appendix B Rotational Invariance via Schur-Weyl
In this section, we demonstrate how certain irreducible representations of $S_{n}$ could correspond to rotationally invariant encoding states when using the encoding specified in Section 4.3.2. Following the work and explanations of Kirby_2018 ; ragone2023representation and in particular the thesis kirby_thesis we reproduce an overview of Schur-Weyl duality and the concept of the Schur basis and add discussion at points linking this to how certain irreducible representation subspaces would correspond to rotationally invariant point cloud encodings in the case considered in Equation 104.
@@ B.1 Introduction to Schur Transforms
The rotational invariant point cloud encoding we propose in Section 4.3.2 relies on a property of the probabilistic irreducible subspace projection circuit when considering the group $S_{n}$, which in effect performs a projection to basis states of the Schur basis. Circuits that implement quantum Schur transforms have previously been introduced, such as those that implement them through the unitary group by iteratively applying the Clebsch-Gordan transforms bacon_efficient_2006 and alternative methods that implement it by considering the action of the symmetric group using quantum Fourier transforms krovi_efficient_2019 . Indeed, further investigation on whether the use of these existing algorithms could be adapted to more efficiently refine our procedure would be an interesting further research direction, as our implementation derives from the most general case of Theorem 4 which applies to any group $G$ and therefore may not be the most efficient in practice.
@@ B.2 Rotational Invariance for Two Points is Trivial
A natural way we choose to represent rotations of points in a quantum encoding is as rotations on the Bloch sphere representation of a qubit as was detailed in Section 4.3.2. This is further motivated by the fact that $SU(2)$ is the double cover of $SO(3)$, which means that every point in $SO(3)$ corresponds to two points in $SU(2)$, which means that we can use $SU(2)$ to model rotations in three-dimensional space classified by $SO(3)$. To produce an encoding like this, we simply need to apply the parameterised rotation gates $R_{x}(\theta)$ and $R_{z}(\phi)$ to an initial $\lvert 0\rangle$ state, where $(\theta,\phi)$ are the angle coordinates of a point in radial coordinates. We are now left with the angular direction of the point represented in the Bloch sphere of a single qubit. If we do this for all $n$ points in the point cloud, we will have $n$ qubits. A rotation of the entire point cloud can now be represented by the application of any $SU(2)$ unitary to every qubit simultaneously. This is a tensor representation of the $SU(2)$ group and can be written explicitly as $\bigotimes_{i=1}^{n}U$ for $U\in\mathrm{SU}(2)$.
To further study this, we shall resort to the case consisting of only two points $n=2$. We again follow standard explanations from reference texts Kirby_2018 ; ragone2023representation ; kirby_thesis while relating these points back to the rotationally invariant encoding we propose throughout the discussion. In this case rotation of the point cloud corresponds to applying the representation $U\otimes U;U\in SU(2)$, to the two qubits. Note that this construction commutes with the SWAP operator $[U\otimes U,\mathrm{SWAP}]=0$, where SWAP corresponds to permuting the two qubits. This fact means that any eigenvectors of $U\otimes U$ with eigenvalue $\lambda$ must also be an eigenvector of SWAP with eigenvalue $\lambda$.
@@
where $\rvert a\rvert^{2}+\rvert b\rvert^{2}=1$. If one applies a Schur transform to $U\otimes U$ which transforms it into the basis described previously it has been shown that it takes a block diagonal form kirby_thesis ; Kirby_2018
@@
The symmetric subspace consists of the upper left block. $U\otimes U$ will transform the state, but keep it within the symmetric subspace. This also means that any variational circuit of the form $U\otimes U$ will similarly keep the state symmetric. This is the intuition behind geometric QML techniques in which variational layers can be constructed, such as $U\otimes U$ in this case, that preserve the symmetry of the state they act on while still allowing variational parameters to be adjusted meyer22 . Although in this work we do not consider creating variational circuits that preserve symmetry, instead we focus on projecting quantum encoded states to irreducible subspaces, which strictly enforces the symmetry invariance in the model as all information in the quantum state is deleted except the portion lying in the chosen irreducible subspace before the training and classification step are considered. It is important to note that the multiplicity of the subspace of $S_{n}$ corresponds to the dimension of the representation of $U\otimes U$ in that subspace and vice versa.

### v2

@@ V Conclusion
The implementation of quantum ResNet demonstrates a procedure that can avoid the trainability issues associated with barren plateaus by allowing shallow depth components to survive in the final loss function. It is still a question of further research as to understanding the exact dynamics of the non-unitary terms. We demonstrated that by parameterising the strength of the residual connections through $\beta_{l}$ terms, one can increase the lower bound probability of success of the LCU method. This provides a possible avenue to tackle the key problem of the LCU method, which is that the implementations can only be performed probabilistically.
We also demonstrated an implementation of quantum native average pooling layers commonly applied in convolutional neural networks. These layers have demonstrated success in classical techniques and therefore have the potential to improve quantum convolutional neural networks. While classically one would need to calculate $\mathcal{O}(N^{2})$ averages, the quantum parallelism inherent in our technique applies the averaging to all pixels (which are amplitude encoded into a quantum state) simultaneously. By generalising our work further, we can recreate quantum convolutional filters as has been previously proposed Wei et al. (2022), while utilising exponentially fewer controlled unitary operations.
Finally, we demonstrated an integration of irreducible representation projections into a quantum framework, allowing for the encoding of data symmetries directly into the quantum state, which has the potential to enhance the model’s generalisation capabilities across various data structures. This general irreducible subspace projection framework includes previous work regarding permutation invariant encodings for point clouds Heredge et al. (2024) as a special case, while in Section IV.3.2 we introduced a novel rotationally invariant encoding for point cloud data using the $S_{4}$ group. We further introduced in Section IV.4 a method of parameterising the amount of symmetry in an encoding as a tuneable hyperparameter. This allows the quantum state to remain in an exponentially large space, while certain subspaces are amplified relative to each other, helping to mitigate some of the simultability concerns associated with fully projecting to a polynomially sized subspace. This added flexibility resulted in an improved performance for point cloud data when encoding an intermediate amount of permutation symmetry when compared to either a fully permutation invariant or noninvariant encoding. These were illustrative examples of how the framework could be utilised and there could be many possible symmetries and datasets for which this framework could be similarly adapted and optimised.
@@
J. Heredge envisaged the project, initially focused on irreducible subspace projections, with M. Sevior and L. Hollenberg as an extension of their previous work Heredge et al. (2024). J. Heredge devised the ResNet and Average Pooling layers applications. The theorems and numerical simulations were developed by J. Heredge with the help of M. West. M. Sevior and L. Hollenberg supervised the overall project. All authors contributed in reviewing the manuscript.
@@ Appendix A Conjugacy Class Implementation
In the implementation of irreducible subspace projections detailed in Section IV, an operator $\hat{\chi}$ was required that contained the character of every element in the group $G$. This meant that the dimension of the matrix scaled with $\rvert G\rvert$, which can be problematic for some groups, such as the permutation group $S_{n}$ where $\rvert S_{n}\rvert=n!$. To tackle this, in this section we will consider a unitary operator $\tilde{\chi}$ that contains the character for each conjugacy class in the group, with additional ancilla qubit registers per conjugacy class prepared in an equal superposition that then acts to distribute these characters to the corresponding group elements. This effectively only requires encoding the character table of the group into a unitary operator, and hence $\tilde{\chi}$ scales with the number of conjugacy classes of $G$. This has the potential to drastically reduce the size of the unitary that must be applied, at the expense of requiring additional ancilla qubits. We show that utilising this circuit structure recreates the result of Theorem 4 and is therefore a valid alternative approach.
The character orthogonality theorem Huppert (1998) stated previously in Equation 91 relates to a sum over all the elements of the group. As the character is the same for all elements in a conjugacy class, this can be adjusted to a sum over all conjugacy classes $\nu_{i}\in\mathcal{C}$ by introducing a term $d_{\nu_{i}}=\rvert\nu_{i}\rvert$ that represents the number of group elements in the conjugacy class $\nu_{i}$. Therefore, the character orthogonality theorem can be written
$\frac{1}{\rvert G\rvert}\sum_{{\nu_{i}}\in\mathcal{C}}d_{\nu_{i}}\chi_{j}({\nu_{i}})\chi_{k}({\nu_{i}})^{*}=\delta_{jk}.$ (117)
@@
$\tilde{\chi}\lvert b_{r}\rangle=\frac{1}{\sqrt{\rvert G\rvert}}\sum_{{\nu_{i}}\in\mathcal{C}}\sqrt{d_{\nu_{i}}}\chi_{r}({\nu_{i}})^{*}\lvert{\nu_{i}}\rangle,$ (118)
@@
This successfully allows for the construction of a unitary matrix $\tilde{\chi}$ that scales in size with the number of conjugacy classes as opposed to the number of group elements. We therefore have satisfied the constraint that $\tilde{\chi}$ is unitary and can be implemented on a quantum device. In general $\tilde{\chi}$ can be written
$\tilde{\chi}=\frac{1}{\sqrt{\rvert G\rvert}}\sum_{r=1}^{R}\sum_{\nu_{i}\in\mathcal{C}}\sqrt{d_{\nu_{i}}}\chi_{r}({\nu_{i}})^{*}\lvert{\nu_{i}}\rangle\langle b_{r}\rvert+\sum_{r=R+1}^{2^{k}}(...)\langle b_{r}\rvert,$ (120)
where ()(...)( … ) collects terms associated with $\lvert b_{r}\rangle,r>R$ which will not be used in the construction.
@@
where $\lvert l_{\nu_{i}}\rangle$ are basis states $\{\lvert l_{\nu_{i}}\rangle\}\in(\mathbb{C}^{2})^{\otimes k_{\nu_{i}}}$ of the $2^{k_{\nu_{i}}}$ dimensional Hilbert space for the $k_{\nu_{i}}$ qubit ancilla register $\lvert\omega_{\nu_{i}}\rangle$, which can be taken to be the computational basis states. The groups elements $l_{\nu_{i}}$ that are contained in the conjugacy class $\nu_{i}$ now label the basis states in register $\lvert\omega_{\nu_{i}}\rangle$. Each register has the requirement $k_{\nu_{i}}\geq\log(d_{\nu_{i}})$ to ensure that there are sufficient qubits in the register such that each element in ${\nu_{i}}$ has a corresponding basis state $\lvert l_{\nu_{i}}\rangle$. The $\lvert\omega_{\nu_{i}}\rangle$ are prepared on a $k_{\nu_{i}}$ qubit register which is initially in a basis state $\lvert b_{1}^{\nu_{i}}\rangle$ through the operation
$\hat{\omega}_{\nu_{i}}=\frac{1}{\sqrt{d_{\nu_{i}}}}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}\lvert l_{\nu_{i}}\rangle\langle b_{1}^{\nu_{i}}\rvert+\sum_{j}(...)\langle b_{j}^{\nu_{i}}\rvert,$ (124)
where ()(...)( … ) summarises terms associated with $\langle b_{j}^{\nu_{i}}\rvert,j\geq 2$ which will not be present in any calculations.
@@
$\otimes\sum_{r=1}^{R}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}a_{r}n_{r}\chi_{r}({\nu_{i}})^{*}\lvert l_{\nu_{i}}\rangle\lvert{\nu_{i}}\rangle U_{{\nu_{i}},l_{\nu_{i}}}\lvert\psi\rangle\Big{)}$
$=\sum_{\nu_{i}}\bigotimes_{{\nu_{i}}^{\prime}\neq{\nu_{i}}}\lvert\omega_{{\nu_{i}}^{\prime}}\rangle\sum_{r=1}^{R}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}\frac{a_{r}n_{r}\chi_{r}({\nu_{i}})^{*}}{\Omega\sqrt{\rvert G\rvert}}\lvert l_{\nu_{i}}\rangle\lvert{\nu_{i}}\rangle$
@@
where we are being flexible with the precise positioning in the tensor product of the unused registers $\lvert\omega_{{\nu_{i}}^{\prime}\neq{\nu_{i}}}\rangle$ for each ${\nu_{i}}^{\prime}$ term for the sake of readability. The notation we are using is that for a given $\nu_{i}$ element in the sum over all conjugacy classes, all $\lvert\omega_{\nu_{i}^{\prime}}\rangle$ registers which will not affect the $S_{\text{SELECT}}$ operator because ${\nu_{i}}^{\prime}\neq{\nu_{i}}$ are written first, with the case ${\nu_{i}}^{\prime}={\nu_{i}}$ written to the right of this. In reality the register position will be different for each $\nu_{i}$, but this becomes difficult to denote in the notation.
We can now uncompute all $\hat{\omega}_{\nu_{i}}$ registers by noting that
$\hat{\omega}_{\nu_{i}}^{\dagger}=\frac{1}{\sqrt{d_{\nu_{i}}}}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}\lvert b_{1}^{\nu_{i}}\rangle\langle l_{\nu_{i}}\rvert+\sum_{j}(...)\lvert b_{j}^{\nu_{i}}\rangle.$ (128)
This means that if we require these registers to be measured in the $\lvert b_{1}^{\nu_{i}}\rangle$ then we can ignore the terms summarised by ()(...)( … ) associated with $\lvert b_{j}^{\nu_{i}}\rangle,j\geq 2$. Application of $\hat{\omega}_{\nu_{i}}^{\dagger}$ to all such registers gives
$\bigotimes_{{\nu_{i}}}\hat{\omega}_{\nu_{i}}^{\dagger}\hat{S}\big{(}\bigotimes_{{\nu_{i}}^{\prime}}\lvert\omega_{{\nu_{i}}^{\prime}}\rangle\otimes\tilde{\chi}\hat{\gamma}\lvert b_{1}\rangle\otimes\lvert\psi\rangle\big{)}$
$=\sum_{\nu_{i}}\bigotimes_{{\nu_{i}}}\lvert b_{1}^{\nu_{i}}\rangle\sum_{r=1}^{R}\sum_{l_{\nu_{i}}=1}^{d_{\nu_{i}}}\frac{a_{r}n_{r}\chi_{r}({\nu_{i}})^{*}}{\Omega\sqrt{d_{\nu_{i}}\rvert G\rvert}}\lvert{\nu_{i}}\rangle\otimes U_{{\nu_{i}},l_{\nu_{i}}}\lvert\psi\rangle$
$+\sum_{j\geq 2}(...)\lvert b_{j}^{\nu_{i}}\rangle$ (129)
We measure the registers which previously contained $\hat{\omega}_{\nu_{i}}$ states and discard unless we measure all $\lvert b_{1}^{\nu_{i}}\rangle$ states $\forall\nu_{i}$. This discards the terms collected by the ()(...)( … ) in the previous equation.
@@
$\langle b_{1}\rvert\tilde{\chi}^{\dagger}=\frac{1}{\sqrt{\rvert G\rvert}}\sum_{{\nu_{i}}\in\mathcal{C}}\sqrt{d_{\nu_{i}}}\chi_{1}({\nu_{i}})\langle{\nu_{i}}\rvert.$ (130)
@@
$\Big{(}\langle b_{1}\rvert\tilde{\chi}^{\dagger}\Big{)}\cdot\Big{(}\bigotimes_{{\nu_{i}}}\langle b_{1}^{\nu_{i}}\rvert\hat{\omega}_{\nu_{i}}^{\dagger}\hat{S}\big{(}\bigotimes_{{\nu_{i}}^{\prime}}\lvert\omega_{{\nu_{i}}^{\prime}}\rangle\otimes\tilde{\chi}\hat{\gamma}\lvert b_{1}\rangle\otimes\lvert\psi\rangle\big{)}\Big{)}$
@@ Appendix B Rotational Invariance via Schur-Weyl
In this section, we demonstrate how certain irreducible representations of $S_{n}$ could correspond to rotationally invariant encoding states when using the encoding specified in Section IV.3.2. Following the work and explanations of Kirby and Strauch (2018); Ragone et al. (2023) and in particular the thesis Kirby we reproduce an overview of Schur-Weyl duality and the concept of the Schur basis and add discussion at points linking this to how certain irreducible representation subspaces would correspond to rotationally invariant point cloud encodings in the case considered in Equation 109.
@@ B.1 Introduction to Schur Transforms
The rotational invariant point cloud encoding we propose in Section IV.3.2 relies on a property of the probabilistic irreducible subspace projection circuit when considering the group $S_{n}$, which in effect performs a projection to basis states of the Schur basis. Circuits that implement quantum Schur transforms have previously been introduced, such as those that implement them through the unitary group by iteratively applying the Clebsch-Gordan transforms Bacon et al. (2006) and alternative methods that implement it by considering the action of the symmetric group using quantum Fourier transforms Krovi (2019). Indeed, further investigation on whether the use of these existing algorithms could be adapted to more efficiently refine our procedure would be an interesting further research direction, as our implementation derives from the most general case of Theorem 4 which applies to any group $G$ and therefore may not be the most efficient in practice.
@@ B.2 Rotational Invariance for Two Points is Trivial
A natural way we choose to represent rotations of points in a quantum encoding is as rotations on the Bloch sphere representation of a qubit as was detailed in Section IV.3.2. This is further motivated by the fact that $SU(2)$ is the double cover of $SO(3)$, which means that every point in $SO(3)$ corresponds to two points in $SU(2)$, which means that we can use $SU(2)$ to model rotations in three-dimensional space classified by $SO(3)$. To produce an encoding like this, we simply need to apply the parameterised rotation gates $R_{x}(\theta)$ and $R_{z}(\phi)$ to an initial $\lvert 0\rangle$ state, where $(\theta,\phi)$ are the angle coordinates of a point in radial coordinates. We are now left with the angular direction of the point represented in the Bloch sphere of a single qubit. If we do this for all $n$ points in the point cloud, we will have $n$ qubits. A rotation of the entire point cloud can now be represented by the application of any $SU(2)$ unitary to every qubit simultaneously. This is a tensor representation of the $SU(2)$ group and can be written explicitly as $\bigotimes_{i=1}^{n}U$ for $U\in\mathrm{SU}(2)$.
To further study this, we shall resort to the case consisting of only two points $n=2$. We again follow standard explanations from reference texts Kirby and Strauch (2018); Ragone et al. (2023); Kirby while relating these points back to the rotationally invariant encoding we propose throughout the discussion. In this case rotation of the point cloud corresponds to applying the representation $U\otimes U;U\in SU(2)$, to the two qubits. Note that this construction commutes with the SWAP operator $[U\otimes U,\mathrm{SWAP}]=0$, where SWAP corresponds to permuting the two qubits. This fact means that $U\otimes U$ and SWAP can be simultaneously diagonalised in a common basis.
@@
where $\rvert a\rvert^{2}+\rvert b\rvert^{2}=1$. If one applies a Schur transform to $U\otimes U$ which transforms it into the basis described previously it has been shown that it takes a block diagonal form Kirby ; Kirby and Strauch (2018)
@@
The symmetric subspace consists of the upper left block. $U\otimes U$ will transform the state, but keep it within the symmetric subspace. This also means that any variational circuit of the form $U\otimes U$ will similarly keep the state symmetric. This is the intuition behind geometric QML techniques in which variational layers can be constructed, such as $U\otimes U$ in this case, that preserve the symmetry of the state they act on while still allowing variational parameters to be adjusted Meyer et al. (2022). Although in this work we do not consider creating variational circuits that preserve symmetry, instead we focus on projecting quantum encoded states to irreducible subspaces, which strictly enforces the symmetry invariance in the model as all information in the quantum state is deleted except the portion lying in the chosen irreducible subspace before the training and classification step are considered. It is important to note that the multiplicity of the subspace of $S_{n}$ corresponds to the dimension of the representation of $U\otimes U$ in that subspace and vice versa.